Action Quest Games Inc. adjusts its accounts annually. The following information is available for the year ended December 31, 2017 1. Purchased a 1-year insurance policy on June 1 for $1,920 cash. 2. Paid $5,600 on August 31 for 5 months rent in advance 3. On September 4, received $3,960 cash in advance from a corporation to sponsor a game each month for a total of 9 months for the most improved students at a local school. 4. Signed a contract for cleaning services starting December 1 for $800 per month. Paid for the first 2 months on November 30. (Hint: Use the account Prepaid Cleaning to record prepayments.) 5. On December 5, received $1,400 in advance from a gaming club. Determined that on December 31, $450 of these games had not yet been played. ▼ (a) Your answer is correct For each of the above transactions, prepare the journal entry to record the initial transaction. (If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when the amount is entered. Do not indent manually.) No. Date Account Titles and Explanation Debit Credit une 1, 2017Prepaid Insurance 1920 Cash 1920 2. Aug. 31, 2017 Prepaid Rent 5600 Cash 5600 TSept. 4, 2017 [cash 3960 Unearned Service Rever 3960 Nov. 30, 2017 Prepaid Cleaning 1600 Cash 1600 Dec. 5, 2017 Cash ▼ 1400 Uneamed Service Rever 1400

Part B Debit/Credit
Answer #1
date account title & explanations debit ($) credit ($) Calculations
Dec-31 Insurance expense 1120 1920*(7/12) = $1120
Prepaid insurance 1120
insuarnce expense for 7 months recorded)
Dec-31 Rent expense 4480 5600*(4/5)= $4480
Prepaid rent 4480
(rent expense for 4 months recorded)
Dec-31 Unearned service revenue 1760 3960*(4/9) = 1760
Service Revenue 1760
(service revenue for 4 months recorded)
Dec-31 Clearing expense 400
Prepaid clearing 400
clearing expense for one month starting from December recorded)
Dec-31 Unearned service revenue 950 1400-450 = 950
Service Revenue 950
(service revenue for $950 recorded)